# Branching Flows: Discrete, Continuous, and Manifold Flow Matching with Splits and Deletions

### v1

@@ Abstract
Diffusion and flow matching approaches to generative modeling have shown promise in domains where the state space is continuous, such as image generation or protein folding & design, and discrete, exemplified by diffusion large language models. They offer a natural fit when the number of elements in a state is fixed in advance (e.g. images), but require ad hoc solutions when, for example, the length of a response from a large language model, or the number of amino acids in a protein chain is not known a priori.
Here we propose Branching Flows, a generative modeling framework that, like diffusion and flow matching approaches, transports a simple distribution to the data distribution. But in Branching Flows, the elements in the state evolve over a forest of binary trees, branching and dying stochastically with rates that are learned by the model. This allows the model to control, during generation, the number of elements in the sequence. We also show that Branching Flows can compose with any flow matching base process on discrete sets, continuous Euclidean spaces, smooth manifolds, and ‘multimodal’ product spaces that mix these components. We demonstrate this in three domains: small molecule generation (multimodal), antibody sequence generation (discrete), and protein backbone generation (multimodal), and show that Branching Flows is a capable distribution learner with a stable learning objective, and that it enables new capabilities.
@@ 1 Introduction
For many applications that involve generating a sequence of elements, it isn’t possible to know the required number of elements in advance. For sequences of discrete tokens this is relatively well addressed by left-to-right autoregressive (AR) sequence models which can control the output length by sampling stop tokens. Diffusion and flow matching methods move beyond left-to-right AR structure, but sacrifice the ability to naturally handle variable-length outputs. Edit Flows, which models variable-length discrete sequences via insertions and deletions, was recently developed (Havasi2025EditFF) and expanded (OneFlowNguyen) to address this.
However for variable-length sequences over continuous spaces there are far fewer options. Continuous autoregressive models are possible (Billera2024TheCL). But in the continuous domain there are practical limitations specifically related to their autoregressive nature. Diffusion and flow models excel for structured grids and fixed-length sequences, but do not naturally handle variable-length outputs. Limited attempts to develop variable-length diffusion models exist (Campbell2023TransDimensionalGM), but these rely on an approximation related to the addition of new elements, and it is unclear how they scale to more complex spaces than the ones upon which they were demonstrated.
@@
One challenge not naturally addressed by fixed-length diffusion/flow models nor AR models (especially in the continuous state context) is what we call the ‘unknown-length infix sampling’ problem, where a segment of unknown length must be inserted into a sequence, conditioned on a priori known flanking regions. Even outside of this context, there is a ‘soft’ version of this problem that arises for fixed-length diffusion/flow models, where flanking regions happen to resolve early during an inference trajectory, and the model cannot control the number of elements spanning them.
@@
Here we present Branching Flows, which is a framework for generative modeling over variable-length sequences, where the elements can be continuous, manifold-valued, discrete, or combinations of these. This is formulated in Generator Matching (Holderrieth2024GeneratorMG), where a conditional process transports samples from a simple distribution to samples from the data distribution. The generator of this conditional process is learned, in expectation, yielding a generator whose marginals match those of the conditional process. Branching Flows augments a base Markov generator on an element space with a branching and deletion process. A forest of trees that represents the branching structure is used to couple a sample from a simple distribution (with elements at the roots) with the data distribution (elements at the leaves). Much like the processes used in molecular evolution and phylogenetics (felsenstein1981evolutionary), the elements evolve independently along each branch but duplicate and decouple at bifurcations in the tree. We show how to use this construction, composed with optional deletions, to train a generative model with naturally hierarchical stochastic paths over sequences of varying length.
We demonstrate Branching Flows on small molecules, antibody sequences, as well as on protein structure generation, which is a real-world multimodal application involving Euclidean, manifold, and discrete states. Finally, we provide a flexible implementation that makes it simple to compose base processes into Branching Flows.
@@ 2 Preliminaries
2.1 Generator Matching
Following Generator Matching (Holderrieth2024GeneratorMG; Lipman2024FlowMG), for $t\in[0,1]$, we seek a probability path $p_{t}$ that interpolates between two fixed marginals $p$ and $q$, where $p$ is typically a distribution that is easy to sample from, and $q$ is the data distribution. This goes via a conditional-path construction: Conditional on an auxiliary variable $z\in\mathcal{Z}$ with marginal distribution $p_{Z}$ and a family of conditional probability measures $p_{t\mid Z}(\mathrm{d}x\mid z)$. The corresponding marginal path can be described by hierarchical sampling
@@ 2.1 Generator Matching
Choose the conditional family so that the boundary constraints $p_{0}=p$ and $p_{1}=q$ hold after marginalization. This is typically done by sampling an $X_{0}\sim p$ and an $X_{1}\sim q$ and having the process start at $X_{0}$ and end at $X_{1}$.
@@
The goal is then to parameterize a generator $\mathcal{L}_{t}$ by a neural network $\mathcal{L}_{t}^{\theta}$ and train it to generate the marginal path $p_{t}$. This is done by first drawing samples from the conditional path $X_{t}\sim p_{t\mid Z}(\mathrm{d}x\mid z)$, and then training $\mathcal{L}_{t}^{\theta}$ against the generators of the conditional paths $\mathcal{L}_{t}^{z}$. Taking gradient steps under a Bregman divergence loss encourages $\mathcal{L}_{t}^{\theta}$ to learn the expectation of the conditional generator which, via a gradient identity, also trains it to generate the marginal path $p_{t}$.
2.2 Augmented Generator Matching
We wish to train a model to sample from a target random variable $X_{1}\sim q(dx)$. By ‘Augmented Generator Matching’ we mean: introduce an auxiliary process $G_{t}$ (with instances denoted by $g_{t}$) for each $t\in[0,1]$, yielding joint training samples $(X_{1},G_{1})\sim p_{1}(dg|x_{1})q(dx)$. Throughout the generative process, the augmentation $G_{t}$ evolves with $X_{t}$. We train a model to emit a generator to transport $(X_{0},G_{0})$ to $(X_{1},G_{1})$, and discard $G_{1}$ to recover samples from the marginal distribution $q(dx)$.
2.3 Counting Flows and Deletion Flows
We first describe Counting Flows and Deletion Flows, which we will use to construct Branching Flows. See Appendix A for a full description of the infinitesimal generator of the process corresponding to having insertions and deletions with these rates, independently for each component.
2.3.1 Poisson Process ‘Counting Flow’
We first introduce ‘Counting Flows’: flow matching over a Poisson counting process that generates samples from the space of non-negative integer ‘counts’ via a conditional process that repeatedly increments by 1 to terminate at an observed count $z_{c}$ by $t=1$. This will later be repurposed for Branching Flows, but we present it here in isolation.
Let $H_{\mathrm{inc}}$ be some hazard distribution supported on $[0,1]$, with its associated hazard rate
$h_{\mathrm{inc}}(t)=\frac{f_{H_{\mathrm{inc}}}(t)}{1-F_{H_{\mathrm{inc}}}(t)}.$
Note that since $H_{\mathrm{inc}}$ is supported on $[0,1]$, we have that $h_{\mathrm{inc}}(t)\to\infty$ as $t\nearrow 1$.
$Q_{t}^{z_{c}}(x_{t}+1|x_{t})$ $=(x_{1}-x_{t})h_{\mathrm{inc}}(t)I_{\{x_{t}<z_{c}\}},$
and $Q_{t}^{z_{c}}(x_{t}|x_{t})=-Q_{t}(x_{t}+1|x_{t}),$ with $Q_{t}^{z_{c}}(\cdot|x_{t})=0$ otherwise.
For any $t$, the increment rate is thus proportional to the number of increments remaining. Due to the exploding rate $Q_{t}^{z_{c}}(x_{t}+1|x_{t})$ as $t\nearrow 1$ when $z_{c}>x_{t}$, the process will almost surely terminate at $z_{c}$.
$D_{R}(R_{t}^{z_{c}},R_{t}^{\theta})=R^{\theta}_{t}-R^{z_{c}}_{t}\log R^{\theta}_{t}$
is a valid loss under an appropriately scaled time-dependent linear parametrization (BilleraTimedep).
Under this process, the distribution of event times is $H$ (see Appendix A.1.1) and for a given $t,x_{t},z_{c}$ we can efficiently sample the waiting time until the next increment event (see Appendix A.1.2) which we will need.
2.3.2 Deletion Flow Matching
We also describe a CTMC in the state space $S=\{0,1\}$, where 0 is ‘present’ and 1 is ‘deleted’. Let $H_{\mathrm{del}}$ be a hazard distribution supported on $[0,1]$, with its associated hazard rate
$h_{\mathrm{del}}(t)=\frac{f_{H_{\mathrm{del}}}(t)}{1-F_{H_{\mathrm{del}}}(t)}$
implying that $h_{\mathrm{del}}(t)\to\infty$ as $t\nearrow 1$. We represent the probability of a deletion by $t=1$ with $Z_{d}\in\{0,1\}$. This is a conditioning variable which determines the conditional rate $Q_{t}^{Z_{d}}(x|x_{t})$, defined by:
$Q^{Z_{d}}_{t}(1|0)=Z_{d}h(t)\quad\text{and}\quad Q_{t}^{Z_{d}}(0|0)=-Q^{Z_{d}}_{t}(1|0),$
along with $Q^{Z_{d}}_{t}(\cdot|x_{t})=0$ for all other values of $x_{t}$. Conditional on $Z_{d}=1$, corresponding to a deleted element, then $Q_{t}^{Z_{d}}(x|x_{t})\to\infty$ as $t\nearrow 1$. Hence, the process $X_{t}|(Z_{d}=1)$ will terminate in deletion at $t=1$ with probability $1$. Let $\rho_{t}^{z_{d}}=z_{d}$ and $\rho_{t}^{\theta}$ be a neural network’s prediction of the probability of deletion by time $t$. Then
is a valid, per-term, conditional generator matching loss under an appropriately scaled time-dependent linear parameterization (BilleraTimedep).
@@ 3 Branching Flows
In Generator Matching, conditioning on $Z$ ensures that the conditional process terminates at a sample from the data distribution, and can be used to control the properties of the conditional paths. For example, ‘minibatch optimal transport’ flow matching (tong2024improvinggeneralizingflowbasedgenerative) draws $X_{0}$ and $X_{1}$ jointly, attempting to minimize the distance between $x_{0},x_{1}$ pairs. With Branching Flows, we make extensive use of $Z$.
@@ 3.1 $Z$
We sample an ordered list of labeled binary rooted plane trees, $\mathcal{T}=(\mathcal{T}_{1},\dots,\mathcal{T}_{L_{0}})$, where the root of each tree $\mathcal{T}_{i}$ is associated with the element $X_{0}^{(i)}$. Each element of $X^{+ø}_{1}$ is associated with a single leaf of a single tree, such that the ordering of the leaves (since the trees are planar) matches the ordering of $X^{+ø}_{1}$. All nodes in the tree are labeled ‘surviving’ except leaf nodes associated with ‘to be deleted‘ states, which are labeled ‘deleted’.
The number of descendants of node $n_{i}$ (including both surviving and deleted) is denoted $w_{i}$. We further decorate the tree with ‘anchors’ $\mathcal{A}$ such that each node has an anchor $(A_{1},\dots)$, where $A_{i}$ is a random variable on $\mathcal{E}$, requiring that the anchor of the surviving leaves equals the leaf-associated element of $X^{+ø}_{1}$ (condition 2), but the internal anchors $\mathcal{A^{I}}$ have no such constraint.
For Branching Flows, $Z\sim p_{Z}$ is a tuple $(X^{+ø}_{1},X_{0},\mathcal{T},\mathcal{A})$, which is obtained by first sampling $X_{1}\sim q$, and then sampling from $p(x^{+ø}_{1},x_{0},\mathcal{T},\mathcal{A}\mid X_{1}=x_{1})$ according to a scheme that satisfies conditions 1 and 2 above, and that also admits a simple marginal $p(x_{0})$. Note that, since $\mathcal{T}$ carries labels for which leaves are deleted, $X_{1}$ is recoverable from $Z$.
To sample $Z$ in practice, to be able to ensure that conditions 1 and 2 are met, we factorize $p(x^{+ø}_{1},x_{0},\mathcal{T},\mathcal{A}\mid x_{1})p(x_{1})$ as follows:
  • Data: $X_{1}\sim q$
  • Initial: $X_{0}\sim\mathcal{D}_{0}|(X_{1}=x_{1})$
  • Data with deletions: $X_{1}^{+ø}\sim\mathcal{D}_{ø}|(X_{1}=x_{1},X_{0}=x_{0})$
  • Forest: $\mathcal{T}\sim\mathcal{D}_{\mathcal{T}}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0})$
  • Internal anchors: $\mathcal{A^{I}}\sim\mathcal{D}|(X_{1}^{+ø}=x_{1}^{+ø},X_{0}=x_{0},\mathcal{T})$ (note: the leaf anchors are already determined by $x_{1}^{+ø}$)
  • $Z:=(X_{1}^{+ø},X_{0},\mathcal{T},\mathcal{A})$
Much of the flexibility in controlling the nature of the flow trajectories comes from flexibility in choice of the above distributions.
@@ 3.2 Conditional Paths
We then construct conditional paths where elements of the state space evolve along branches of our trees. To do this, each element must track which branch on which tree it belongs to, and so we augment the state space with a branch indicator $g_{t}$. We thus have $\tilde{X}_{t}:=(X_{t},G_{t})\sim p_{t|Z}(\mathrm{d}(x,g)\mid z)$ that begins at $\tilde{x}_{0}:=(x_{0},g_{0})$ and terminates at $\tilde{x}_{1}:=(x_{1},g_{1})$, where at any $t$ each element can split into two, or be deleted, depending on the trees $\mathcal{T}$ in $Z$.
and analogously define $\tilde{x}_{t}^{(i)}\triangleright\downarrow$.
$\operatorname{split}_{i}((\tilde{x}_{t}^{(i)})):=(\tilde{x}_{t}^{(1)},\ldots,(\tilde{x}_{t}^{(i)}\triangleright\uparrow),\ (\tilde{x}_{t}^{(i)}\triangleright\downarrow),\ldots,\tilde{x}_{t}^{(n)}).$
that deletes the element $\tilde{x}^{(i)}$ at index $i$.
Between split events, each element evolves along the branch it is associated with independently of the other elements according to the chosen ‘base’ generator but conditioned to terminate, were it to reach $t=1$, at the anchor $a_{i}$ of the node ahead of it.
For each element, the instantaneous split event rate is just as in our counting flow described above, where a split hazard distribution $H_{\operatorname{split}}$ is chosen, but with $w_{i}-1$ remaining splits to be realized by $t=1$. The split event intensity is thus $h_{\operatorname{split}}(t)\cdot(w_{i}-1)$. Similarly, a deletion hazard distribution $H_{\operatorname{del}}$ is chosen, and the deletion event intensity is $h_{\operatorname{del}}(t)$ on branches associated with deleted leaves, but 0 on branches associated with internal and surviving terminal nodes.
Since the hazard rates are designed such that all splits and all deletions implied by $\mathcal{T}$ will be consumed by $t=1$, and since the anchors $a_{i}$ of all surviving terminal nodes are equal to the elements of $x_{1}$, we thus have a conditional path $x_{t}\sim p_{t\mid Z}(x_{t}\mid z)$ that begins at $x_{0}$ and terminates at $x_{1}$ with probability $1$.
See figure 2 for a depiction of $Z$ and the conditional path construction. Let $\tilde{\mathcal{E}}$ denote the augmented component-wise state space, and recall that each component-wise state space is simply $\mathcal{E}$. In the future, e.g., when describing conditional paths and the loss, we will denote $\tilde{x}_{t}\in\bigsqcup_{n\geq 1}\tilde{\mathcal{E}}^{n}$ and $\tilde{X}_{t}$ for a $\bigsqcup_{n\geq 1}\tilde{\mathcal{E}}^{n}$ valued random variable in the main text, while in the appendix we may omit the decoration for brevity of notation.

### v3

@@ Abstract
Diffusion and flow matching approaches to generative modeling have shown promise in domains where the number of elements in a state is fixed in advance (e.g. images), but require ad hoc solutions when, for example, the length of a response from a large language model, the number of atoms in a molecule, or the number of amino acids in a protein chain is not known a priori. Here we propose Branching Flows, a generative modeling framework that, like diffusion and flow matching approaches, transports a simple distribution to the data distribution. But in Branching Flows, the elements in the state evolve over a forest of binary trees, branching and dying stochastically with rates that are learned by the model. This allows the model to control, during generation, the number of elements in the sequence. We show that Branching Flows can compose with any flow matching base process on discrete sets, continuous Euclidean spaces, Riemannian manifolds, and ‘multimodal’ product spaces that mix these components, and we demonstrate distribution matching on small molecules and antibody sequences, and that this scales to complicated domains such as protein structures.
@@ 1 Introduction
For many applications that involve generating a sequence of elements, it isn’t possible to know the required number of elements in advance. For sequences of discrete tokens this is well addressed by autoregressive (AR) models which control length by sampling stop tokens. Continuous AR models are possible (Billera2024TheCL) but underexplored, and current examples are limited in the range of problems they can address. Diffusion and flow matching methods avoid left-to-right AR structure and excel for structured grids and fixed-length sequences, but sacrifice the ability to naturally handle length variation.
@@
One challenge not naturally addressed by fixed-length diffusion/flow models nor AR models (especially in the continuous state context) is what we call the ‘unknown-length infix sampling’ problem, where a segment of unknown length must be inserted into a sequence, conditioned on a priori known flanking regions. As a concrete example from protein design: conditioned on framework regions and some binding pose, if you wish to design the complementarity-determining regions of an antibody you must first specify how long they must be, which is absurd. There is a ‘soft’ version of this problem that arises for fixed-length diffusion/flow models, where flanking regions happen to resolve early during an inference trajectory, and the model cannot control the number of elements spanning them.
In the diffusion and flow matching context, Edit Flows addresses length variation by modeling variable-length discrete sequences via insertions and deletions (Havasi2025EditFF; OneFlowNguyen) and, in the purely-continuous domain, one attempt to develop variable-length diffusion models (Campbell2023TransDimensionalGM) relies on an approximation related to the addition of new elements, and it is unclear how this scales to more complex examples.
@@
Here we present Branching Flows, a framework for generative modeling over variable-length sequences where the elements can be continuous, manifold-valued, discrete, or combinations of these. This is formulated in Generator Matching (Holderrieth2024GeneratorMG), where a conditional process transports samples from a simple distribution to the data distribution. The generator of this is learned, in expectation, yielding a process whose marginals match those of the conditional process. Branching Flows augments a base Markov generator on an element space with a branching and deletion process. A forest of trees is used to couple a sample from a simple distribution (with elements at the roots) with the data distribution (elements at the leaves). Like the processes used in phylogenetics (felsenstein1981evolutionary), the elements evolve independently along each branch but duplicate and decouple at bifurcations in the tree. This construction, composed with optional deletions, allows us to train a generative model over sequences of varying length.
@@ 2 Preliminaries
2.1 Generator Matching and Auxiliary Generator Matching
Following Generator Matching (Holderrieth2024GeneratorMG; Lipman2024FlowMG), for $t\in[0,1]$, we seek a probability path $p_{t}$ that interpolates between two fixed marginals $p$ and $q$, where $p$ is a distribution that is easy to sample from, and $q$ is the data distribution. This goes via a conditional-path construction: Conditional on an auxiliary variable $z\in\mathcal{Z}$ with marginal distribution $p_{Z}$ and a family of conditional probability measures $p_{t\mid Z}(\mathrm{d}x\mid z)$, the corresponding marginal path can be described by hierarchical sampling
@@ 2.1 Generator Matching and Auxiliary Generator Matching
Choose the conditional family so that the boundary constraints $p_{0}=p$ and $p_{1}=q$ hold after marginalization. This can be achieved by sampling an $X_{0}\sim p$ and an $X_{1}\sim q$ and having the process start at $X_{0}$ and end at $X_{1}$. The goal is then to parameterize a generator $\mathcal{L}_{t}$ by a neural network $\mathcal{L}_{t}^{\theta}$ and train it to generate the marginal path $p_{t}$. This is done by first drawing samples from the conditional path $X_{t}\sim p_{t\mid Z}(\mathrm{d}x\mid z)$, and then training $\mathcal{L}_{t}^{\theta}$ against the generators of the conditional paths $\mathcal{L}_{t}^{z}$. Taking gradient steps under a ‘conditional generator matching’ loss (a Bregman divergence) encourages $\mathcal{L}_{t}^{\theta}$ to learn the expectation of the conditional generator which, via a gradient identity, generates the marginal path $p_{t}$.
By ‘Auxiliary Generator Matching’ we mean: augment our process with $G_{t}$ (with realisations denoted $g_{t}$) for each $t\in[0,1]$, yielding joint conditional paths $(X_{t},G_{t})|(Z=z)\sim p_{t\mid Z}(\mathrm{d}x,\mathrm{d}g\mid z)$, but train a model that marginalizes over $G_{t}$ and $Z$, learning a generator on $X_{t}$ with marginals matching $X_{t}\sim p_{t}(\mathrm{d}x)$. This was used in Edit Flows (Havasi2025EditFF) for the discrete case, but see Appendix E.3 for our setting.
Figure 2: Branching Flows construction. Left outlines the sampling of $Z$ when $x_{1}$ has 4 elements and $x_{0}$ has two elements, for a conditional bridge that will incur a single deletion (element $x_{1}^{+(2)ø})$). $m(.,.)$ derives an anchor from two children, which could be e.g. a midpoint for continuous states and a mask token for discrete states. Right depicts the conditional sampling of $x_{t}|Z$, where elements can split and be deleted but only, in the conditional path, according to the pre-determined branching structure of the two trees $\mathcal{T}$ in $Z$.
@@ 3 Branching Flows
In Generator Matching, conditioning on $Z$ ensures that the conditional process terminates at a sample from the data distribution, and can be used to control the properties of the conditional paths (e.g. ‘minibatch optimal transport’ flow matching, tong2024improvinggeneralizingflowbasedgenerative, which samples from coupled $X_{0}$ and $X_{1}$ jointly). With Branching Flows, we make extensive use of $Z$, which we first describe in section 3.1, and then describe the conditional paths given $Z$ in section 3.2.
@@ 3.1 $Z$
To couple $X_{0}$ to $X^{+ø}_{1}$, we sample an ordered list of labeled binary rooted plane trees, $\mathcal{T}=(\mathcal{T}_{1},\dots,\mathcal{T}_{L_{0}})$, where the root of each tree $\mathcal{T}_{i}$ is associated with the element $X_{0}^{(i)}$. Each element of $X^{+ø}_{1}$ is associated with a single leaf of a single tree, such that the ordering of the leaves (since the trees are planar) matches the ordering of $X^{+ø}_{1}$. All nodes in the tree are labeled ‘surviving’ except leaf nodes associated with ‘to be deleted‘ states, which are labeled ‘deleted’. The number of descendants of node $n_{i}$ (including surviving and deleted) is denoted $w_{i}$.
To steer the conditional process along internal branches, we introduce ‘anchors’ $\mathcal{A}$, where each node’s anchor $A_{i}$ is a random variable on $\mathcal{E}$, requiring the anchor of the surviving leaves to equal the leaf-associated element of $X^{+ø}_{1}$ (condition 2). Internal anchors $\mathcal{A^{I}}$ have no such constraint, and can e.g. be chosen to steer the process along internal branches towards the centroid of the descendant elements.
For Branching Flows, $Z\sim p_{Z}$ is a tuple $(X^{+ø}_{1},X_{0},\mathcal{T},\mathcal{A})$, which is obtained by first sampling $X_{1}\sim q$, and then sampling from $p(x^{+ø}_{1},x_{0},\mathcal{T},\mathcal{A}\mid X_{1}=x_{1})$ according to a scheme that satisfies conditions 1 and 2 above, and that also admits a simple marginal $p(x_{0})$. Note that, since $\mathcal{T}$ carries labels for which leaves are deleted, $X_{1}$ is recoverable from $Z$. To sample $Z$ in practice, to be able to ensure that conditions 1 and 2 are met, we factorize $p(x^{+ø}_{1},x_{0},\mathcal{T},\mathcal{A}\mid x_{1})p(x_{1})$ as shown in fig. 2. See App. A for concrete schemes for sampling $Z$, especially $\mathcal{T}$ and $\mathcal{A}$.
@@ 3.2 Conditional Paths
We construct conditional paths where elements evolve along branches of our trees, splitting and deleting, to terminate at $X_{1}$. We need to specify i) how the elements evolve along a branch, ii) what happens when they split, iii) the per-element split rate, and iv) the per-element deletion rate.
Between split or deletion events, each element evolves independently of the other elements along the branch with which it is associated. This is governed by a ‘base’ generator, which can be any element-wise conditional process generator from Generator Matching, allowing drift, diffusion, and jumps, and conditioned to terminate at a specific value by $t=1$. In Branching Flows, the anchor associated with each branch steers the base generator, which is conditioned to terminate, were it to reach $t=1$, at the anchor $a_{i}$ of the node ahead of it.
When an element is split it is replaced by two adjacent duplicates, and each duplicate is associated with a child branch. To control the rates of split events, let $H_{\mathrm{split}}$ be a ‘split hazard’ distribution supported on $[0,1]$ with no atom at $1$, with its associated hazard rate $h_{\mathrm{split}}(t)=\frac{f_{H_{\mathrm{split}}}(t)}{1-F_{H_{\mathrm{split}}}(t)}$ where $f_{H_{\mathrm{split}}}(t)$ is the density at $t$ and $F_{H_{\mathrm{split}}}(t)$ is the cumulative distribution. For an element associated with a specific branch, the tree (which is pre-sampled, within $Z$) determines the number of remaining split events $w_{i}-1$ to be realized by $t=1$. The instantaneous per-element split event rate is $(w_{i}-1)\cdot h_{\operatorname{split}}(t)$, and it holds that all remaining splits are consumed by $t=1$ when $w_{i}>1$.
For deletions, let $H_{\mathrm{del}}$ similarly define $h_{\mathrm{del}}(t)$, and the deletion event rate is $h_{\operatorname{del}}(t)$ on branches associated with deleted leaves, and 0 on branches associated with internal and surviving terminal nodes. Since the hazard rates are designed such that all splits and all deletions implied by $\mathcal{T}$ will be consumed by $t=1$, and since the anchors $a_{i}$ of all surviving terminal nodes are equal to the elements of $x_{1}$, we thus have a conditional path $x_{t}\sim p_{t\mid Z}(x_{t}\mid z)$ that begins at $x_{0}$ and terminates at $x_{1}$ with probability $1$.
For each element to track which branch on which tree it belongs to (which is required for the conditional process to be Markov in the state), we augment the state space with a branch indicator $g_{t}$. With each element in $\tilde{\mathcal{E}}$, we have $\tilde{X}_{t}:=(X_{t},G_{t})\sim p_{t|Z}(\mathrm{d}x,\mathrm{d}g|z)$ that begins at $\tilde{x}_{0}:=(x_{0},g_{0})$ and terminates at $\tilde{x}_{1}:=(x_{1},g_{1})$. See Appendix C for details of our branch tracking and split and deletion operators, and see figure 2 for a depiction of $Z$ and the conditional path construction.